OpenStudy (anonymous):

What is the value of the discriminant for the quadratic equation 2x2 - 3x + 1 = 0? 1 -17 -1 -15

OpenStudy (anonymous):

\[b^2 - 4ac\] \[b^2 - 4ac = (-3)^2 - 4(2)(1)\]

OpenStudy (anonymous):

then continue..

OpenStudy (anonymous):

\[(-3)^2 - 4(2) (1) = 9 - 8 = 1 \]

OpenStudy (anonymous):

Any questions? so you use \[b^2 - 4ac\]

OpenStudy (anonymous):

No, thank you!